# Reusing Historical Trajectories in Natural Policy Gradient via Importance Sampling: Convergence and Convergence Rate

### v1

@@ Abstract
Keywords importance sampling  $\cdot$ reinforcement learning
@@ 1 Introduction
    We provide a rigorous asymptotic convergence analysis of the proposed algorithm by the ODE approach. We further characterize the improved convergence rate by the SDE approach.
@@ 2.1 Preliminaries: Markov Decision Process
The policies can also be parameterized by neural networks, where significant empirical successes have been achieved in many challenging applications, such as playing Go (see [23]). The performance of a policy is evaluated in terms of the expected discounted total return
$\eta(\pi_{\theta})=\mathbb{E}_{s_{0},a_{0},\ldots}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}\left(s_{t},a_{t}\right)\right],$
where $s_{0}\sim\rho_{0}\left(s_{0}\right),a_{t}\sim\pi_{\theta}\left(a_{t}\mid s_{t}\right),s_{t+1}\sim\mathcal{P}\left(s_{t+1}\mid s_{t},a_{t}\right)$. Denote by $d^{\pi_{\theta}}(s)$ the discounted state visitation distribution induced by the policy $\pi_{\theta}$, $d^{\pi_{\theta}}(s)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathcal{P}\left(s_{t}=s\mid\pi_{\theta}\right)$. It is useful to define the discounted occupancy measure as $d^{\pi_{\theta}}(s,a)=d^{\pi_{\theta}}(s)\pi_{\theta}(a|s)$. Using the discounted occupancy measure, we can rewrite the expected discounted return as $\eta(\pi_{\theta})=\mathbb{E}_{(s,a)\sim d^{\pi_{\theta}}(s,a)}[\mathcal{R}(s,a)]$. The goal is for the agent to find the optimal policy $\pi_{\theta^{*}}$ that maximizes the expected discounted return, or equivalently, $\theta^{*}=\arg\max_{\theta\in\Theta}\eta(\pi_{\theta})$. We use the following standard definitions of the value function $V^{\pi_{\theta}}$, the state-action value function $Q^{\pi_{\theta}}$, and the advantage function $A^{\pi_{\theta}}$: $V^{\pi_{\theta}}\left(s_{t}\right)=\mathbb{E}_{a_{t},s_{t+1},\ldots}\left[\sum_{l=0}^{\infty}\gamma^{l}R\left(s_{t+l},a_{t+l}\right)\right]$, $Q^{\pi_{\theta}}\left(s_{t},a_{t}\right)=\mathbb{E}_{s_{t+1},a_{t+1},\ldots}\left[\sum_{l=0}^{\infty}\gamma^{l}R\left(s_{t+l},a_{t+l}\right)\right]$, and $A^{\pi_{\theta}}(s,a)=Q^{\pi_{\theta}}(s,a)-V^{\pi_{\theta}}(s)$.
@@ 2.3 Natural Policy Gradient with Reusing Historical Trajectories
The independence can be then satisfied by generating independent trajectories with random lengths. However, it is worth noting in this way only the last sample is utilized and it requires re-starting the environment for each sample. While the i.i.d. assumption is necessary to demonstrate the convergence of the proposed algorithm, as also assumed in, e.g., [24, 25], in practice, the algorithm is usually implemented in a more sample-efficient way, such as single-path generation in [22]. The empirical efficiency of the proposed algorithm is demonstrated in Section 5, even when i.i.d. data are not available. A vanilla baseline FIM estimator $\widetilde{F}\left(\theta_{n}\right)$ and gradient estimator $\widetilde{\nabla\eta}\left(\theta_{n}\right)$ can be obtained as:
$\widetilde{F}\left(\theta_{n}\right)=\frac{1}{B}\sum_{i=1}^{B}S(\xi_{n}^{i},\theta_{n}),\quad\widetilde{\nabla\eta}\left(\theta_{n}\right)=\frac{1}{B}\sum_{i=1}^{B}G(\xi_{n}^{i},\theta_{n}),$
where $S(\xi,\theta)=\nabla\log\pi_{\theta}(a|s)(\nabla\log\pi_{\theta}(a|s))^{T}$ and $G(\xi,\theta)=\frac{1}{1-\gamma}A^{\pi_{\theta}}(s,a)\nabla\log\pi_{\theta}(a|s)$, where $\xi$ appears in $A^{\pi_{\theta}}(s,a)$ and $\pi_{\theta}(a|s)$. It is easy to see that $\widetilde{F}\left(\theta_{n}\right)$ and $\widetilde{\nabla\eta}\left(\theta_{n}\right)$ are unbiased estimators of the FIM $F(\theta_{n})$ and the gradient $\nabla\eta\left(\theta_{n}\right)$, respectively. However, in the vanilla stochastic natural policy gradient (VNPG), a small batch size $B$, which is often the case when there is limited online interaction with the environment, could lead to a large variance in the estimator. An alternative FIM and gradient estimator, which reuse historical trajectories, are as follows:
$\widehat{F}\left(\theta_{n}\right)=\frac{1}{KB}\sum_{m=n-K+1}^{n}\sum_{i=1}^{B}\omega(\xi_{m}^{i},\theta_{n}|\theta_{m})S(\xi_{m}^{i},\theta_{n}),$
$\widehat{\nabla\eta}(\theta_{n})=\frac{1}{KB}\sum_{m=n-K+1}^{n}\sum_{i=1}^{B}\omega(\xi_{m}^{i},\theta_{n}|\theta_{m})G(\xi_{m}^{i},\theta_{n}),$ (2)
where we reuse previous $K-1$ iterations’ trajectories. The likelihood ratio $\omega(\xi_{m}^{i},\theta_{n}|\theta_{m})$ is given by
@@
$\widehat{F}\left(\theta_{n}\right)=\epsilon I_{d}+\frac{1}{KB}\sum_{m=n-K+1}^{n}\sum_{i=1}^{B}\omega(\xi_{m}^{i},\theta_{n}|\theta_{m})S(\xi_{m}^{i},\theta_{n}).$ (4)
@@
As pointed out by [21] and prior works [19], and [20], the dependence between iterations introduces bias into the FIM and gradient estimators when reusing historical trajectories.
@@
Let’s use a two-iteration example to illustrate this bias. Consider the first step of RNPG given a deterministic initial solution $\theta_{0}\in\Theta$. Note that both the FIM estimator $\widehat{F}(\theta_{0})$ and gradient estimator $\widehat{\nabla\eta}(\theta_{0})$ are based on $B$ replications run at $\theta_{0}$. Therefore, we have $\theta_{1}=\operatorname{Proj}_{\Theta}\left(\theta_{0}+\alpha_{0}(\widehat{F}^{-1}(\theta_{0})\widehat{\nabla\eta}(\theta_{0}))\right)$. For simplicity, assume no projection is needed for $\theta_{1}$, thus $\theta_{1}=\theta_{0}+\alpha_{0}(\widehat{F}^{-1}(\theta_{0})\widehat{\nabla\eta}(\theta_{0}))$. Also note that $\widehat{\nabla\eta}(\theta_{1})=\frac{1}{2B}\sum_{i=1}^{B}G(\xi_{0}^{i},\theta_{1})\omega(\xi_{0}^{i},\theta_{1}|\theta_{0})+\frac{1}{2B}\sum_{i=1}^{B}G(\xi_{1}^{i},\theta_{1})$, $\widehat{F}(\theta_{1})=\frac{1}{2}\left(\epsilon I_{d}+\frac{1}{B}\sum_{i=1}^{B}S(\xi_{0}^{i},\theta_{1})\omega(\xi_{0}^{i},\theta_{1}|\theta_{0})\right)+\frac{1}{2}\left(\epsilon I_{d}+\frac{1}{B}\sum_{i=1}^{B}S(\xi_{1}^{i},\theta_{1})\right)$. Then the expectation of the gradient estimator in RNPG is
@@
$\mathbb{E}[\widehat{\nabla\eta}(\theta_{1})|\theta_{1}]=\frac{1}{2}\mathbb{E}\left[\frac{1}{B}\sum_{i=1}^{B}G(\xi_{0}^{i},\theta_{1})\omega(\xi_{0}^{i},\theta_{1}|\theta_{0})\bigg{|}\theta_{1}\right]+\frac{1}{2}\nabla\eta(\theta_{1}),$
where the first term can be written as
@@
$\frac{1}{2}\mathbb{E}\left[\frac{1}{B}\sum_{i=1}^{B}G(\xi_{0}^{i},\theta_{1})\omega(\xi_{0}^{i},\theta_{1}|\theta_{0})\bigg{|}\theta_{1}=\theta_{0}+\alpha_{0}\left(\epsilon I_{d}+\frac{1}{B}\sum_{i=1}^{B}S(\xi_{0}^{i},\theta_{0})\right)^{-1}\frac{1}{B}\sum_{i=1}^{B}G(\xi_{0}^{i},\theta_{0})\right]$
@@
$=$ $\frac{1}{2}\mathbb{E}\left[\frac{1}{B}\sum_{i=1}^{B}G(\xi_{0}^{i},\theta_{1})\omega(\xi_{o}^{i},\theta_{1}|\theta_{0})\bigg{|}\frac{1}{B}\sum_{i=1}^{B}G(\xi_{0}^{i},\theta_{0})=\left(\epsilon I_{d}+\frac{1}{B}\sum_{i=1}^{B}S(\xi_{0}^{i},\theta_{0})\right)\frac{\theta_{1}-\theta_{0}}{\alpha_{0}}\right]$
$\neq$ $\frac{1}{2}\nabla\eta(\theta_{1}).$
Similarly, the natural policy gradient estimator in RNPG is also biased. In summary, the conditional distribution of $\xi_{0}^{i}$ given $\theta_{1}$ differs from the distribution from which $\xi_{0}^{i}$ was originally sampled. It should also be noted that the likelihood ratio in (4) and (2) is usually hard to compute, since the discounted occupancy measure does not admit a closed form expression. We defer the discussion to Section 3 on some approximations to make Algorithm 1 more practical. The main theoretical result is summarized as follows: (1) the solution trajectory $\{\theta_{n}\}$ in Algorithm 1 converges with probability one (w.p.1) to a limiting point $\bar{\theta}$ such that $\nabla\eta(\bar{\theta})=0$; (2) the normalized error $E_{n}=\frac{\theta_{n}-\bar{\theta}}{\sqrt{\alpha_{n}}}$ converges weakly to a stationary solution to the SDE $dE=\mathcal{G}Edt+dW$, where $W$ is a Wiener process with covariance matrix $\widehat{\Sigma}(\bar{\theta})=\frac{1}{B}\Sigma_{1}(\bar{\theta})+\frac{1}{KB}\Sigma_{2}(\bar{\theta})$. The matrices $\mathcal{G}$, $\Sigma_{1}$, $\Sigma_{2}$ will be explained in Section 4. In essence, the proposed RNPG algorithm maintains the convergence and improves the convergence rate by an order of $O(\frac{1}{K})$.

### v2

@@ Abstract
*These authors contributed equally.
Keywords Importance sampling  $\cdot$ reinforcement learning  $\cdot$ policy gradient
@@ 1 Introduction
    We provide a rigorous asymptotic convergence analysis of the proposed algorithm by an ODE approach. We further characterize the improved convergence rate by an SDE approach.
@@ 2.1 Preliminaries: Markov Decision Process
The policies can also be parameterized by neural networks, where significant empirical successes have been achieved in many challenging applications, such as playing Go (see [27]). The goal is to find an optimal policy $\pi_{\theta^{*}}$ that maximizes the expected total discounted return:
$\theta^{*}\in\arg\max_{\theta\in\Theta}\eta(\pi_{\theta}):=\mathbb{E}_{s_{0},a_{0},\ldots}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}\left(s_{t},a_{t}\right)\right],$ (1)
where $s_{0}\sim\rho_{0}\left(s_{0}\right),a_{t}\sim\pi_{\theta}\left(a_{t}\mid s_{t}\right),s_{t+1}\sim\mathcal{P}\left(s_{t+1}\mid s_{t},a_{t}\right)$, $\Theta$ is the feasible set of $\theta$, and $\eta(\pi_{\theta})$ is the expected total discounted return of policy $\pi_{\theta}$.
Denote by $d^{\pi_{\theta}}(s)$ the discounted state visitation distribution induced by the policy $\pi_{\theta}$, $d^{\pi_{\theta}}(s)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathcal{P}\left(s_{t}=s\mid\pi_{\theta}\right)$. It is useful to define the discounted occupancy measure as $d^{\pi_{\theta}}(s,a)=d^{\pi_{\theta}}(s)\pi_{\theta}(a|s)$. Using the discounted occupancy measure, we can rewrite the expected total discounted return as $\eta(\pi_{\theta})=\mathbb{E}_{(s,a)\sim d^{\pi_{\theta}}(s,a)}[\mathcal{R}(s,a)]$. We use the following standard definitions of the value function $V^{\pi_{\theta}}$, the state-action value function $Q^{\pi_{\theta}}$, and the advantage function $A^{\pi_{\theta}}$: $V^{\pi_{\theta}}\left(s_{t}\right)=\mathbb{E}_{a_{t},s_{t+1},\ldots}\left[\sum_{l=0}^{\infty}\gamma^{l}R\left(s_{t+l},a_{t+l}\right)\right]$, $Q^{\pi_{\theta}}\left(s_{t},a_{t}\right)=\mathbb{E}_{s_{t+1},a_{t+1},\ldots}\left[\sum_{l=0}^{\infty}\gamma^{l}R\left(s_{t+l},a_{t+l}\right)\right]$, and $A^{\pi_{\theta}}(s,a)=Q^{\pi_{\theta}}(s,a)-V^{\pi_{\theta}}(s)$.
@@ 2.3 Natural Policy Gradient with Reusing Historical Trajectories
The independence can be then satisfied by generating independent trajectories with random lengths. However, it is worth noting in this way only the last sample is utilized and it requires re-starting the environment for each sample. While the i.i.d. assumption is necessary to demonstrate the convergence of the proposed algorithm, as also assumed in, e.g., [23, 13], in practice, the algorithm is usually implemented in a more sample-efficient way, such as single-path generation in [25]. The empirical efficiency of the proposed algorithm is demonstrated in Section 5, even when i.i.d. data are not available. A vanilla baseline gradient estimator $\widetilde{\nabla\eta}\left(\theta_{n}\right)$ and FIM estimator $\widetilde{F}\left(\theta_{n}\right)$ can be obtained as:
$\widetilde{\nabla\eta}\left(\theta_{n}\right)=\frac{1}{B}\sum_{i=1}^{B}G(\xi_{n}^{i},\theta_{n}),\quad\widetilde{F}\left(\theta_{n}\right)=\frac{1}{B}\sum_{i=1}^{B}S(\xi_{n}^{i},\theta_{n}),$
where $S(\xi,\theta)=\nabla\log\pi_{\theta}(a|s)(\nabla\log\pi_{\theta}(a|s))^{T}$ and $G(\xi,\theta)=\frac{1}{1-\gamma}A^{\pi_{\theta}}(s,a)\nabla\log\pi_{\theta}(a|s)$, where $\xi$ appears in $A^{\pi_{\theta}}(s,a)$ and $\pi_{\theta}(a|s)$. It is easy to see that $\widetilde{F}\left(\theta_{n}\right)$ and $\widetilde{\nabla\eta}\left(\theta_{n}\right)$ are unbiased estimators of the FIM $F(\theta_{n})$ and the gradient $\nabla\eta\left(\theta_{n}\right)$, respectively. However, in the vanilla stochastic natural policy gradient (VNPG), a small batch size $B$, which is often the case when there is limited online interaction with the environment, could lead to a large variance in the estimator. An alternative gradient and FIM estimator, which reuse historical trajectories, are as follows:
$\widehat{\nabla\eta}(\theta_{n})=\frac{1}{K_{1}B}\sum_{m=n-K_{1}+1}^{n}\sum_{i=1}^{B}\omega(\xi_{m}^{i},\theta_{n}|\theta_{m})G(\xi_{m}^{i},\theta_{n}),$ (3)
$\widehat{F}\left(\theta_{n}\right)=\frac{1}{K_{2}B}\sum_{m=n-K_{2}+1}^{n}\sum_{i=1}^{B}\omega(\xi_{m}^{i},\theta_{n}|\theta_{m})S(\xi_{m}^{i},\theta_{n}),$
where we reuse previous $K_{1}-1$ iterations’ trajectories for estimating the gradient and $K_{2}-1$ iterations’ trajectories for estimating FIM. The likelihood ratio $\omega(\xi_{m}^{i},\theta_{n}|\theta_{m})$ is given by
@@
$\widehat{F}\left(\theta_{n}\right)=\epsilon I_{d}+\frac{1}{K_{2}B}\sum_{m=n-K_{2}+1}^{n}\sum_{i=1}^{B}\omega(\xi_{m}^{i},\theta_{n}|\theta_{m})S(\xi_{m}^{i},\theta_{n}).$ (5)
@@
From a computational perspective, it should also be noted that the likelihood ratio in (5) and (3) is usually hard to compute, since the discounted occupancy measure does not admit a closed form expression. We defer the discussion to Section 3 on some approximations to make Algorithm 1 more practical.
2.4 Bias and Variance of the Gradient Estimator Reusing Samples from Previous Iteration
@@
Some prior work, including [17], [7], and [6], have considered reusing past samples via the likelihood ratio approach in simulation optimization problems and noted that the dependence between iterations introduces bias into their gradient estimators. Similar to their observations, such iteration dependence also introduces bias into our FIM and gradient estimators when reusing historical trajectories. Let’s use a two-iteration example ($K_{1}=K_{2}=K=2$) to illustrate this bias. Consider the first step of RNPG given a deterministic initial solution $\theta_{0}\in\Theta$. Note that both the FIM estimator $\widehat{F}(\theta_{0})$ and gradient estimator $\widehat{\nabla\eta}(\theta_{0})$ are based on $B$ replications run at $\theta_{0}$. Therefore, we have $\theta_{1}=\operatorname{Proj}_{\Theta}\left(\theta_{0}+\alpha_{0}(\widehat{F}^{-1}(\theta_{0})\widehat{\nabla\eta}(\theta_{0}))\right)$. For simplicity, assume no projection is needed for $\theta_{1}$, thus $\theta_{1}=\theta_{0}+\alpha_{0}(\widehat{F}^{-1}(\theta_{0})\widehat{\nabla\eta}(\theta_{0}))$. Also note that $\widehat{\nabla\eta}(\theta_{1})=\frac{1}{2B}\sum_{i=1}^{B}G(\xi_{0}^{i},\theta_{1})\omega(\xi_{0}^{i},\theta_{1}|\theta_{0})+\frac{1}{2B}\sum_{i=1}^{B}G(\xi_{1}^{i},\theta_{1})$, $\widehat{F}(\theta_{1})=\frac{1}{2}\left(\epsilon I_{d}+\frac{1}{B}\sum_{i=1}^{B}S(\xi_{0}^{i},\theta_{1})\omega(\xi_{0}^{i},\theta_{1}|\theta_{0})\right)+\frac{1}{2}\left(\epsilon I_{d}+\frac{1}{B}\sum_{i=1}^{B}S(\xi_{1}^{i},\theta_{1})\right)$. Then the expectation of the gradient estimator in RNPG is
@@
$\mathbb{E}[\widehat{\nabla\eta}(\theta_{1})|\theta_{1}]=\underbrace{\frac{1}{2}\mathbb{E}\left[\frac{1}{B}\sum_{i=1}^{B}G(\xi_{0}^{i},\theta_{1})\omega(\xi_{0}^{i},\theta_{1}|\theta_{0})\bigg{|}\theta_{1}\right]}_{Z}+\frac{1}{2}\nabla\eta(\theta_{1}).$
Since $\theta_{1}=\theta_{0}+\alpha_{0}\left(\epsilon I_{d}+\frac{1}{B}\sum_{i=1}^{B}S(\xi_{0}^{i},\theta_{0})\right)^{-1}\frac{1}{B}\sum_{i=1}^{B}G(\xi_{0}^{i},\theta_{0})$ depends on $\xi_{0}^{i},i=1,2,\ldots,B$, $\xi_{0}^{i}|\theta_{1}$ does not follow the occupancy measure $d^{\pi_{\theta_{0}}}$. This means
$\mathbb{E}\left[G(\xi_{0}^{i},\theta_{1})\omega(\xi_{0}^{i},\theta_{1}|\theta_{0})\bigg{|}\theta_{1}\right]$
$\neq$ $\int_{\xi}G(\xi,\theta_{1})\omega(\xi,\theta_{1}|\theta_{0})d^{\pi_{\theta_{0}}}(\xi)\mathrm{d}\xi$
$=$ $\int_{\xi}G(\xi,\theta_{1})\frac{d^{\pi_{\theta_{1}}}(\xi)}{d^{\pi_{\theta_{0}}}(\xi)}d^{\pi_{\theta_{0}}}(\xi)\mathrm{d}\xi$
$=$ $\int_{\xi}G(\xi,\theta_{1})d^{\pi_{\theta_{1}}}(\xi)\mathrm{d}\xi$
$=$ $\nabla\eta(\theta_{1}).$
@@
Hence, $Z=\frac{1}{2}\sum_{i=1}^{B}\mathbb{E}\left[G(\xi_{0}^{i},\theta_{1})\omega(\xi_{0}^{i},\theta_{1}|\theta_{0})\bigg{|}\theta_{1}\right]\neq\nabla\eta(\theta_{1})$.
By extension, the natural policy gradient estimator in RNPG is also biased. In summary, the conditional distribution of $\xi_{0}^{i}$ given $\theta_{1}$ differs from the distribution from which $\xi_{0}^{i}$ was originally sampled.
In addition to the bias, let’s also consider the variance of the natural gradient estimator when using historical trajectories. Similarly, consider the gradient estimator using samples from the first and second iteration as discussed above. We can write the variance of the gradient estimator (conditioned on $\theta_{1}$) as:
@@
$\operatorname{VaR}\left(\widehat{\nabla\eta}(\theta_{1})\right)=\underbrace{\frac{1}{4}\operatorname{VaR}\left(\frac{1}{B}\sum_{i=1}^{B}G(\xi_{0}^{i},\theta_{1})\omega(\xi_{0}^{i},\theta_{1}|\theta_{0})\bigg{|}\theta_{1}\right)}_{Z^{\prime}}+\frac{1}{4B}\operatorname{VaR}(G(\xi,\theta_{1})),$ (7)
where the covariance term
$2\mathbb{E}\left[\left(\frac{1}{2B}\sum_{i=1}^{B}G(\xi_{0}^{i},\theta_{1})\omega(\xi_{0}^{i},\theta_{1}|\theta_{0})-Z\right)\left(\frac{1}{B}\sum_{i=1}^{2B}G(\xi_{1}^{i},\theta_{1})-\frac{1}{2}\nabla\eta(\theta_{1})\right)\bigg{|}\theta_{1}\right]$
$=$ $2\mathbb{E}\left[\left(\frac{1}{2B}\sum_{i=1}^{B}G(\xi_{0}^{i},\theta_{1})\omega(\xi_{0}^{i},\theta_{1}|\theta_{0})-Z\right)\bigg{|}\theta_{1}\right]\mathbb{E}\left[\left(\frac{1}{B}\sum_{i=1}^{2B}G(\xi_{1}^{i},\theta_{1})-\frac{1}{2}\nabla\eta(\theta_{1})\right)\bigg{|}\theta_{1}\right]$
$=$ $0,$
where the first equality holds since conditioned on $\theta_{1}$, $\xi_{1}^{i}$ is independent of past samples. From (7), compared with the variance of the gradient estimator only using the current samples, which is $\frac{1}{B}\operatorname{VaR}(G(\xi,\theta_{1}))$, reusing samples in $\widehat{\nabla\eta}(\theta_{1})$ will reduce the conditional variance if $Z^{\prime}\leq\frac{3}{4B}\operatorname{VaR}\left(G(\xi,\theta_{1})\right)$ and increase the conditional variance otherwise. The value of $Z^{\prime}$ depends on the specific form of the importance ratio $\omega$ and the random gradient $G$.
Nonetheless, from an asymptotic perspective, reusing past samples is beneficial. Before diving into the rigorous analysis, we provide some intuitive explanations here. First, as discussed above, the bias introduced by reusing samples is due to the dependence of current solution on past samples. With a diminishing step size sequence $\{\alpha_{n}\}$, the difference between solutions $\theta_{n}$ and $\theta_{n+\ell}$ with a fixed $\ell$ vanishes as $n$ goes to infinity. This indicates $\theta_{n+\ell}\approx\theta_{n}$ and $\omega(\cdot,\theta_{n+\ell}|\theta_{n})\approx 1$ when $n$ is large, and as a result, asymptotically, we can view all samples generated between iterations $n$ and $n+\ell$ as approximately conditional i.i.d. samples under parameter $\theta_{n+\ell}\approx\theta_{n}$. This further indicates the bias caused by the difference among distributions of samples from iteration $n$ to $n+\ell$ also decreases to 00 asymptotically. Similarly for the variance, as reused samples become asymptotically i.i.d., we expect the conditional variance of the gradient estimator decreases as the reuse size $K$ increases. In summary, we expect by reusing samples, the bias is negligible and the variance is reduced asymptotically. Throughout the remaining paper, we will show this rigorously and demonstrate numerically.
2.5 Summary of Main Theoretical Results
We first give an intuitive summary of the main theoretical results in this paper, and an interested reader can go on to the next two sections for the detailed analysis. Choosing $K_{1}=K,K_{2}=1$ (i.e., using samples of the most recent $K$ iterations in the gradient estimator, and only using samples of the current iteration in the FIM estimator), we show the following results:
  • (1)
    the solution trajectory $\{\theta_{n}\}$ given by RNPG (Algorithm 1) converges with probability one (w.p.1) to a local optimum $\bar{\theta}$, which is the same as the vanilla natural policy gradient (VNPG).
  • (2)
    the random error $\theta_{n}-\bar{\theta}$ follows a normal distribution $\mathcal{N}(0,\alpha_{n}\Sigma_{\infty})$ asymptotically, where $\alpha_{n}$ is the step size and $\Sigma_{\infty}\propto\frac{1}{B}\Sigma_{1}(\bar{\theta})+\frac{1}{KB}\Sigma_{2}(\bar{\theta})$. The definition of $\Sigma_{\infty}$, $\Sigma_{1}$ and $\Sigma_{2}$ will be explained in Section 4. In essence, $\sqrt{\alpha_{n}}A$, where $A$ satisfies $AA^{T}=\Sigma_{\infty}$, characterizes the rate of convergence of $\theta_{n}$. We show that the proposed RNPG algorithm improves the convergence rate via decreasing the magnitude of $\Sigma_{2}$ by a factor of $\frac{1}{K}$.